A)Use the Keynesian model to calculate equilibrium income values, assuming the price level is fixed.

C=375 +0.75(Y-T) I=575 G=200 T=200 X = 700 M= 100 +0.15 Y

Equilibrium Income (Y)=

B) Analyze the effect of expansionary fiscal policy in this open economy model. Specifically, assume the government raises spending and lowers taxes to 180; calculate the new value of equilibrium income

New Equilibrium Income (Y)=
Answer #1

At equilibrium, Y = 375 + 0.75(Y-200) + 575 + 200 + 700 - 100 - 0.15Y

Y = 1600 + 0.6Y

0.4Y = 1600

Y= 4000

B) government spending multiplier = 1/(1-0.75+0.15) = 2.5

Tax spending multiplier = -0.75/(1-0.75+0.15) = 1.875

Change in equilibrium income = 2.5*180 - 1.875(-180) = 787.5

New equilibrium income = 4000 + 787.5= 4787.5